# Ultrasonic Medical Tissue Imaging Using Probabilistic Inversion: Leveraging Variational Inference for Speed Reconstruction and Uncertainty Quantification

### v2

@@ Abstract
Full Waveform Inversion (FWI) is a promising technique for achieving high-resolution imaging in medical ultrasound. Traditional FWI methods suffer from issues related to computational efficiency, dependence on initial models, and the inability to quantify uncertainty. This study introduces the Stein Variational Gradient Descent (SVGD) algorithm into FWI, aiming to improve inversion performance and enhance uncertainty quantification. By deriving the posterior gradient, the study explores the integration of SVGD with FWI and demonstrates its ability to approximate complex priors. In-silico experiments with synthetic data and real-world breast tissue data highlight the advantages of the SVGD-based framework over conventional FWI. SVGD-based FWI improves inversion quality, provides more reliable uncertainty quantification, and offers a tighter bound for the prior distribution. These findings show that probabilistic inversion is a promising tool for addressing the limitations of traditional FWI methods in ultrasonic imaging of medical tissues.
@@ 1 Introduction
Traditional FWI algorithms face several challenges, including low computational efficiency, strong dependence on initial models, and difficulties in accurately quantifying parameter uncertainty [16]. In uncertainty estimation for FWI, Markov Chain Monte Carlo (MCMC) methods, such as the Metropolis-Hastings algorithm [17], have traditionally been used to sample from complex posterior distributions [18, 19]. However, MCMC methods are inefficient in high-dimensional spaces due to their random-walk behavior, which leads to slow convergence and the tendency to become trapped at individual maxima in the probability distribution [20]. Although more advanced MCMC techniques, such as reversible-jump MCMC [21, 22, 23], stochastic Newton MCMC[24, 25], and Hamiltonian Monte Carlo [26, 27], aim to improve efficiency, these methods remain intractable for FWI due to their extremely high computational cost.
@@
To improve the efficiency of MCMC methods, Variational Inference (VI) techniques are increasingly applied to estimate the posterior distribution of model parameters and address uncertainties inherent in the inversion process [28, 29, 30]. Instead of directly calculating the posterior distribution, VI approximates it by minimizing the Kullback-Leibler (KL) divergence between the true posterior and a simpler, tractable distribution [31, 32]. Recent research, such as in stochastic variational inference (SVI) for medical ultrasound imaging [33], has demonstrated how gradient-based methods can efficiently solve inverse problems. The derivation of SVI depends on selecting a suitable prior distribution, which can be challenging to determine in practice. Stein Variational Gradient Descent (SVGD), a VI-based method, uses a set of particles to approximate more complex prior distributions, such as a Gaussian mixture model [34]. The SVGD approach enhances both the efficiency and accuracy of VI in high-dimensional inverse problems, making it a promising method for addressing challenges in ultrasound imaging.
@@
This paper explores the application of SVGD in medical ultrasound imaging, focusing on enhancing the efficiency and robustness of the inversion process. The core principles of the wave equation used in FWI to model wave propagation are outlined, followed by a discussion of the challenges in model inversion due to noise and complex data. The paper then introduces VI as a method for incorporating uncertainty into the inversion process, improving upon traditional optimization techniques. SVGD is presented as an advanced VI-based algorithm that optimizes the approximation of posterior distributions. The gradient of the posterior distribution is derived, and SVGD is applied to solve inverse problems in medical ultrasound imaging. Through simulations using both synthetic and real-world medical data, the results demonstrate that SVGD leads to faster convergence, improves image reconstruction quality, and provides more reliable uncertainty quantification.
@@
This paper presents a theoretical framework for applying the SVGD algorithm to FWI. Section 2 introduces the foundational concepts of FWI and VI, providing the necessary background to understand their integration within a probabilistic framework. Section 3 focuses on the application of SVGD to FWI, including the derivation of the posterior gradient and an example demonstrating SVGD’s ability to approximate Gaussian Mixture Model. In Section 4, the framework is applied to synthetic data generated from linear-array and ring-array transducers. Section 5 extends the methodology to breast tissue data. Both Sections 4 and 5 use conventional FWI as a baseline to evaluate the inversion quality of SVGD. Section 5 also employs the Stochastic Variational Inference (SVI) for a comparative analysis of the uncertainty quantification provided by SVGD.
2 Probabilistic Inversion in FWI: Bayesian Framework and Variational Approximations
@@ 2.1 Acoustic Wave Propagation and Data Generation in FWI
The wave equation in Full Waveform Inversion (FWI) is used to simulate wave propagation and iteratively update the model to match the observed data. The most commonly used wave equation is the acoustic wave equation [35, 36],
$\frac{1}{\bm{c}^{2}}\frac{\partial^{2}\bm{u}(\bm{x},t)}{\partial t^{2}}-\nabla^{2}\bm{u}(\bm{x},t)=\bm{s}(\bm{x},t),$ (1)
where $u(\bm{x},t)$ represents the wavefield, such as pressure or displacement, at position $\bm{x}=(x_{1},x_{2},\cdots,x_{d})$ and time $t$. The notation $\bm{c}=\bm{c}(\bm{x})$ is the speed of sound (SOS), which varies spatially in a heterogeneous medium. The operator $\nabla^{2}=\sum_{i}\partial^{2}/\partial x_{i}^{2}$ is the Laplacian operator, representing the spatial derivatives of the wavefield. The source signal $\bm{s}(\bm{x},t)$ describes the spatial and temporal distribution of the energy input into the system. The exact form of $\bm{s}(\bm{x},t)$ depends on the specific modeling of the ultrasound source. For a simple point source, the expression is
@@
The receivers, positioned at $\bm{x}_{r}$, convert the mechanical vibrations of the incoming acoustic waves into electrical signals. These signals consist of both reflected and transmitted waves. To predict the wavefield data recorded by the receivers, the wave equation (1) is numerically solved using the finite difference method. The obtained data $\bm{d}_{r}$ is a time series and described by a functional relation that involves the position $\bm{x}_{r}$, time $t$, and material properties $\bm{m}$ of the medium. This relationship is
@@ 2.2 Misfit Function and Optimization in FWI
FWI begins with an initial model of the medium, which serves as an approximation of the model parameters $\bm{m}$ at each spatial point. Using this initial model, the predicted wavefield data $\bm{d}_{r}$ is computed numerically by solving the wave equation (1). The core of FWI is the iterative updating of the parameters $\bm{m}$ by minimizing the difference between $\bm{d}_{r}^{\text{obs}}$ and $\bm{d}_{r}$. The difference between the two datasets is quantified using a misfit function, often based on the least-squares error. The most commonly used misfit function is the $l_{2}$ norm, which measures the sum of squared differences across all receivers,
@@
The goal of FWI is to find the optimal parameters $\bm{m}^{*}$ that minimize the misfit function, i.e., to solve the following optimization problem:
@@
2.3 Variational Inference Solves Bayes’ Equation
In recent years, Variational Inference (VI) has emerged as a powerful tool for improving the efficiency and robustness of inversion techniques. VI provides an alternative approach to classical optimization by approximating complex posterior distributions with simpler, tractable distributions, thus reducing the computational burden. The goal of FWI is to find the optimal parameters $\bm{m}^{*}$ satisfying Eq. (5), VI treats the parameters $\bm{m}$ as random variables and infers their distribution given the observed data. This allows for more reliable reconstructions of tissue properties, providing both the best estimate of model parameters and a measure of uncertainty that is crucial for clinical decision-making. By approximating the true posterior distribution through a variational family, VI facilitates efficient model updating and allows for more robust handling of noisy or incomplete data. This is beneficial in medical ultrasound imaging, where measurement noise, sensor limitations, and complex tissue structures often make traditional deterministic inversion methods prone to local minima or overfitting. VI-based methods can enhance the accuracy and stability of FWI, providing more reliable reconstructions of tissue properties.
@@ 2.3 Variational Inference Solves Bayes’ Equation
where $p(\bm{m})$ is the prior distribution that encodes our initial beliefs about the model parameters before observing the data. The evidence distribution $p(\bm{d}^{\text{obs}})$ (also known as marginal likelihood) acts as a normalizing constant ensuring that the posterior distribution integrates to one. The likelihood distribution $p(\bm{d}^{\text{obs}}|\bm{m})$ represents the probability of observing the data given a set of model parameters.
Computing the exact posterior distribution in Eq. (6) often requires solving complex integrals, which can be computationally intractable. To illustrate this, consider the joint distribution between the prior and the likelihood, i.e., $p(\bm{m},\bm{d}^{\text{obs}})=p(\bm{d}^{\text{obs}}|\bm{m})p(\bm{m})$. The evidence $p(\bm{d}^{\text{obs}})$ is the marginal of the joint distribution $p(\bm{m},\bm{d}^{\text{obs}})$,
[…]
The numerical integrate the integral in Eq. (7) is often challenging due to the high dimensionality of the model parameters $\bm{m}$. As a result, directly computing the posterior becomes intractable.
Another widely used numerical method to solve the posterior distribution in Eq. (6) is Markov Chain Monte Carlo (MCMC) [17, 37]. MCMC is a class of algorithms that generates samples from a probability distribution, allowing for an approximation of the posterior distribution. MCMC constructs a Markov chain, and the distribution of the chain’s states approaches the target posterior distribution after a large number of iterations. The Monte Carlo estimate of the posterior distribution $\hat{p}(\bm{m}|\bm{d}^{\text{obs}})$ is obtained by averaging over the samples generated by the chain,
$\hat{p}(\bm{m}|\bm{d}^{\text{obs}})=\frac{1}{c}\sum_{i}p(\bm{d}^{\text{obs}}|\bm{m}^{i})p(\bm{m}^{i}),$ (8)
where $p(\bm{m}^{i})$ represents the prior distribution evaluated at the $i$-th sample. The notation $c$ is the normalizing constant ensuring that the estimated posterior integrates to one. Although MCMC is a useful tool, it is computationally intensive, requiring numerous iterations to achieve accurate results. Besides, MCMC may struggle with high-dimensional spaces or complex dependencies between parameters.
The posterior distribution $p(\bm{m}|\bm{d}^{\text{obs}})$ is often computationally challenging to obtain using analytical integral or numerical MCMC, which limits their applicability in FWI. To address these limitations, VI approximates the posterior distribution $p(\bm{m}|\bm{d}^{\text{obs}})$ by minimizing the KL divergence between the true posterior and a simpler, tractable variational distribution $q(\bm{m})$. The variational distribution $q(\bm{m})$ belongs to a chosen family of distributions $\mathcal{Q}$, and the optimal distribution $q^{*}(\bm{m})$ is obtained by
@@
[…]
@@
[…]
$D_{KL}\left[q(\bm{m})\|p(\bm{m}|\bm{d}^{\text{obs}})\right]=E_{q}\left[\log(q(\bm{m}))\right]-E_{q}\left[\log(p(\bm{m}|\bm{d}^{\text{obs}}))\right].$ (11)
Substituting Bayes’ theorem (6) into Eq. (11),
$D_{KL}\left[q(\bm{m})\|p(\bm{m}|\bm{d}^{\text{obs}})\right]=E_{q}\left[\log(q(\bm{m}))\right]-E_{q}\left[\log(p(\bm{m},\bm{d}^{\text{obs}}))\right]+E_{q}\left[\log(p(\bm{d}^{\text{obs}}))\right].$ (12)
In the above equation, the evidence $\log(p(\bm{d}^{\text{obs}}))$ is difficult to compute directly. Consequently, minimizing the KL divergence cannot be done directly. Rearranging the Eq. (12) gives
$E_{q}\left[\log(p(\bm{m},\bm{d}^{\text{obs}}))\right]-E_{q}\left[\log(q(\bm{m}))\right]=\log(p(\bm{d}^{\text{obs}}))-D_{KL}\left[q(\bm{m})\|p(\bm{m}|\bm{d}^{\text{obs}})\right].$ (13)
Since $D_{\text{KL}}\left[q(\bm{m})\|p(\bm{m}|\bm{d}^{\text{obs}})\right]\geq 0$, a lower bound for the evidence is given by
$\log(p(\bm{d}^{\text{obs}}))\geq E_{q}\left[\log(p(\bm{m},\bm{d}^{\text{obs}}))\right]-E_{q}\left[\log(q(\bm{m}))\right].$ (14)
The right-hand defines the Evidence Lower Bound (ELBO), denoted by
$\text{ELBO}(q)\triangleq E_{q}\left[\log(p(\bm{m},\bm{d}^{\text{obs}}))\right]-E_{q}\left[\log(q(\bm{m}))\right].$ (15)
Substituting the definition of ELBO into Eq. (13) yields
$\text{ELBO}(q)=\log(p(\bm{d}^{\text{obs}}))-D_{KL}\left[q(\bm{m})\|p(\bm{m}|\bm{d}^{\text{obs}})\right].$ (16)
Because $\log(p(\bm{d}^{\text{obs}}))$ is constant for a specific problem, minimizing the KL divergence in Eq. (9) is equivalent to maximizing the ELBO. The optimization problem in Eq. (9) becomes
$q^{*}(\bm{m})=\mathop{\arg\max}_{q(\bm{m})\in\mathcal{Q}}\text{ELBO}(q).$ (17)
The definition of $\text{ELBO}(q)$ does not involve the intractable term $p(\bm{d}^{\text{obs}})$, maximizing the ELBO is computationally easier than directly minimizing the KL divergence. Observing Eq. (14) and (16),
$\log(p(\bm{d}^{\text{obs}}))=\text{ELBO}(q)\ \text{if and only if}\ D_{\text{KL}}\left[q(\bm{m})\|p(\bm{m}|\bm{d}^{\text{obs}})\right]=0.$
The maximal value of $\text{ELBO}(q)$ is achieved when the approximating distribution $q(\bm{m})$ is exactly equal to the posterior distribution $p(\bm{m}|\bm{d}^{\text{obs}})$.
By deriving the ELBO, VI reduces the computational burden compared to exact Bayesian inference methods, such as analytically solving Eq. (7) or using MCMC methods. The flexibility of VI lies in the choice of optimization techniques and the variational family $\mathcal{Q}$, which leads to the development of various algorithms. Among these, Stein Variational Gradient Descent (SVGD) stands out as a flexible and efficient method. In the following, this paper introduces SVGD and demonstrates its efficiency in solving inverse problems.
3 SVGD-Based Posterior Approximation and Uncertainty Quantification in FWI
3.1 Overview of SVGD

### v3

@@ Abstract
Background and Objective: Full Waveform Inversion (FWI) is a promising technique for achieving high-resolution imaging in medical ultrasound. However, conventional FWI methods suffer from issues related to computational efficiency, dependence on initial models, and the inability to quantify uncertainty. This study aims to enhance inversion performance and provide a reliable method for uncertainty quantification in medical FWI imaging.
Methods: This study integrates the Stein Variational Gradient Descent (SVGD) algorithm into the FWI framework by deriving the posterior gradient for probabilistic inversion. To evaluate the proposed method, numerical experiments are conducted on synthetic datasets, including a breast tissue model with realistic anatomical structure. Imaging accuracy and uncertainty quantification are assessed to compare the performance of SVGD-based FWI with conventional FWI and Stochastic Variational Inference (SVI) methods. Markov Chain Monte Carlo (MCMC) is implemented as a benchmark to evaluate the quality of uncertainty estimates.
Results: For synthetic data, the SVGD-based FWI framework yields more precise estimates in the region of interest (ROI) and demonstrates faster convergence compared to the conventional FWI. For the anatomically realistic breast tissue simulation, SVGD produces a maximum relative error of 1.10% and a mean relative error of 0.09% in the ROI. The estimated uncertainty is spatially consistent, with most values below 0.01 and a mean of approximately 0.003. Compared to SVI, SVGD provides improved structural resolution and stronger agreement with the MCMC benchmark, indicating more reliable uncertainty quantification.
Conclusions: The SVGD-based FWI method improves inversion quality, enhances uncertainty quantification. These findings indicate that probabilistic inversion is a promising tool for addressing the limitations of traditional FWI methods in ultrasonic imaging of medical tissues.
@@ 1 Introduction
Traditional FWI algorithms face several challenges, including low computational efficiency, strong dependence on initial models, and difficulties in accurately quantifying parameter uncertainty [16]. In uncertainty estimation for FWI, Markov Chain Monte Carlo (MCMC) methods, such as the Metropolis-Hastings algorithm [17], have traditionally been used to sample from complex posterior distributions [18, 19]. However, MCMC methods are inefficient in high-dimensional spaces due to their random-walk behavior, which leads to slow convergence and the tendency to become trapped at individual maxima in the probability distribution [20, 21]. Although more advanced MCMC techniques, such as reversible-jump MCMC [22, 23, 24], stochastic Newton MCMC[25, 26], and Hamiltonian Monte Carlo [27, 28], aim to improve efficiency, these methods remain intractable for FWI due to their extremely high computational cost.
@@
To improve the efficiency of MCMC methods, Variational Inference (VI) techniques are increasingly applied to estimate the posterior distribution of model parameters and address uncertainties inherent in the inversion process [29, 30, 31]. Instead of directly calculating the posterior distribution, VI approximates it by minimizing the Kullback-Leibler (KL) divergence between the true posterior and a simpler, tractable distribution [32, 33]. Recent research, such as stochastic variational inference (SVI) for medical ultrasound imaging [34], has demonstrated how gradient-based methods can efficiently solve inverse problems. The derivation of SVI depends on selecting a suitable prior distribution, which can be challenging to determine in practice. Stein Variational Gradient Descent (SVGD), a VI-based method, uses a set of particles to approximate more complex prior distributions, such as a Gaussian mixture model [35]. The SVGD approach enhances both the efficiency and accuracy of VI in high-dimensional inverse problems, making it a promising method for addressing challenges in ultrasound imaging.
@@
This paper explores the application of SVGD in medical ultrasound imaging, focusing on enhancing the efficiency and robustness of the inversion process. The core principles of the wave equation used in FWI to model wave propagation are outlined, followed by a discussion of the challenges in model inversion due to noise and complex data. The paper then introduces VI as a method for incorporating uncertainty into the inversion process, improving upon traditional optimization techniques. SVGD is presented as an advanced VI-based algorithm that optimizes the approximation of posterior distributions. The gradient of the posterior distribution is derived, and SVGD is applied to solve inverse problems in medical ultrasound imaging. Simulation results on both synthetic phantoms and anatomically realistic breast models demonstrate that SVGD improves convergence speed, reconstruction accuracy, and the reliability of uncertainty estimates.
@@
The structure of this paper is outlined below. Section 2 introduces the foundational concepts of FWI and VI, with a focus on the application of SVGD to FWI, including the derivation of the posterior gradient. In Section 3, the proposed framework is applied to synthetic data generated from both linear-array and ring-array transducers, and further extended to breast tissue data. This section uses conventional FWI as a baseline to evaluate the inversion quality of SVGD, and employs SVI and MCMC for a comparative analysis of the uncertainty quantification provided by SVGD. Finally, Sections 4 and 5 provide a discussion of the results and conclude the study, respectively.
2 Methods
@@ 2.1 Acoustic wave propagation and data generation in FWI
The wave equation in FWI is used to simulate wave propagation and iteratively update the model to match the observed data. The most commonly used wave equation is the acoustic wave equation [36, 37],
$\frac{1}{\bm{c}^{2}}\frac{\partial^{2}\bm{p}(\bm{x},t)}{\partial t^{2}}-\nabla^{2}\bm{p}(\bm{x},t)=\bm{s}(\bm{x},t),$ (1)
where $\bm{p}(\bm{x},t)$ represents acoustic pressure at position $\bm{x}=(x_{1},x_{2},\cdots,x_{d})$ and time $t$. The notation $\bm{c}=\bm{c}(\bm{x})$ is the speed of sound (SOS), which varies spatially in a heterogeneous medium. In FWI applications, the model parameter is often reparameterized as $\bm{m}=1/\bm{c}^{2}$, known as the squared slowness, to facilitate gradient-based optimization and improve numerical convergence. The operator $\nabla^{2}=\sum_{i}\partial^{2}/\partial x_{i}^{2}$ is the Laplacian operator, representing the spatial derivatives of the pressure field. The source signal $\bm{s}(\bm{x},t)$ describes the spatial and temporal distribution of the energy input into the system. The exact form of $\bm{s}(\bm{x},t)$ depends on the specific modeling of the ultrasound source. For a simple point source, the expression is
@@
The receivers, positioned at $\bm{x}_{r}$, convert the mechanical vibrations of the incoming acoustic waves into electrical signals. These signals consist of both reflected and transmitted waves. To predict the wavefield data recorded by the receivers, the wave equation (1) is numerically solved using the finite difference method. The obtained data $\bm{d}_{r}$ is a time series and described by a functional relation that involves the position $\bm{x}_{r}$, time $t$, and squared slowness $\bm{m}$. This relationship is
@@
FWI begins with an initial model of the medium, which serves as an approximation of the model parameters $\bm{m}$ at each spatial point. Using this initial model, the predicted wavefield data $\bm{d}_{r}$ is computed numerically by solving the wave equation (1). The core of FWI is the iterative updating of the parameters $\bm{m}$ by minimizing the difference between $\bm{d}_{r}^{\text{obs}}$ and $\bm{d}_{r}$. The difference between the two datasets is quantified using a misfit function, often based on the least-squares error. The most commonly used misfit function is the $l_{2}$ norm, which measures the sum of squared differences across all receivers [34, 38, 39],
@@
The goal of FWI is to find the optimal parameters $\bm{m}^{*}$ that minimize the misfit function, i.e., to solve the following optimization problem,
@@
2.2 Variational inference solves FWI problems
In recent years, VI has emerged as a powerful tool for improving the efficiency and robustness of inversion techniques. VI provides an alternative approach to classical optimization by approximating complex posterior distributions with simpler, tractable distributions, thus reducing the computational burden. The goal of FWI is to find the optimal parameters $\bm{m}^{*}$ satisfying Eq. (5), VI treats the parameters $\bm{m}$ as random variables and infers their distribution given the observed data. This allows for more reliable reconstructions of tissue properties, providing both the best estimate of model parameters and a measure of uncertainty that is crucial for clinical decision-making. By approximating the true posterior distribution through a variational family, VI facilitates efficient model updating and allows for more robust handling of noisy or incomplete data. This is beneficial in medical ultrasound imaging, where measurement noise, sensor limitations, and complex tissue structures often make traditional deterministic inversion methods prone to local minima or overfitting. VI-based methods can enhance the accuracy and stability of FWI, providing more reliable reconstructions of tissue properties.
@@ 2.2 Variational inference solves FWI problems
where $p(\bm{m})$ is the prior distribution that encodes our initial beliefs about the model parameters before observing the data. The evidence distribution $p(\bm{d}^{\text{obs}})$ acts as a normalizing constant ensuring that the posterior distribution integrates to one. The likelihood distribution $p(\bm{d}^{\text{obs}}|\bm{m})$ represents the probability of observing the data given a set of model parameters. Although Eq.(4) does not explicitly include a regularization term, such terms are incorporated in practice to mitigate ill-posedness and improve inversion stability. In the Bayesian framework, regularization naturally arises through the prior distribution, where the negative logarithm of the prior acts as a regularizer. This probabilistic interpretation provides a principled way to incorporate prior knowledge into the inversion.
A widely used numerical method to approximate the posterior distribution in Eq. (6) is MCMC [17, 40]. MCMC is a class of algorithms that generates samples from a distribution proportional to the posterior, i.e., $p(\bm{m}|\bm{d}^{\text{obs}})\propto p(\bm{d}^{\text{obs}}|\bm{m})p(\bm{m})$. The samples allow for estimating expectations and quantifying uncertainty without explicitly computing the normalization constant.
Although MCMC provides an asymptotically correct approximation of the posterior, it can be computationally intensive and inefficient in high-dimensional inverse problems such as FWI. To address these limitations, VI approximates the posterior distribution $p(\bm{m}|\bm{d}^{\text{obs}})$ by minimizing the KL divergence between the true posterior and a simpler, tractable variational distribution $q(\bm{m})$. The variational distribution $q(\bm{m})$ belongs to a chosen family of distributions $\mathcal{Q}$, and the optimal distribution $q^{*}(\bm{m})$ is obtained by
@@
$D_{KL}\left[q(\bm{m})\|p(\bm{m}|\bm{d}^{\text{obs}})\right]=$ $\int q(\bm{m})\log\left(\frac{q(\bm{m})}{p(\bm{m}|\bm{d}^{\text{obs}})}\right)d\bm{m}$
$=$ $\int q(\bm{m})\log\left(q(\bm{m})\right)d\bm{m}-\int q(\bm{m})\log\left(p(\bm{m}|\bm{d}^{\text{obs}})\right)d\bm{m}.$ (8)
@@
By leveraging the KL divergence, VI reduces the computational burden compared to exact Bayesian inference methods, such as MCMC methods. The flexibility of VI lies in the choice of optimization techniques and the variational family $\mathcal{Q}$, which leads to the development of various algorithms. Among these, SVGD stands out as a flexible and efficient method. In the following, this paper introduces SVGD and demonstrates its efficiency in solving inverse problems.
2.3 SVGD-based posterior approximation and uncertainty quantification in FWI